# GENIE Collaboration

### v1

@@ IV.2 Alternative Tune: G24-t
A similar improvement is achieved by the intermediate tune, G24-t, illustrated in Fig. 12. Instead of significantly increasing proton PIPD to elevate the $p_{\textrm{N}}$ tail, this tune notably enhances $R_{\textrm{SRC}}$—less-energetic final states are present and therefore an increased RES/DIS (Fig. 12a) ratio. Rather than suppressing $S_{\textrm{CEX}}^{\pi}$ (cf. Fig. 16d in Appendix .2), this tune significantly increases it (Fig. 12b) to offset the further reduction of “No Interaction” $\pi^{0}$ events (smaller $S_{\lambda}^{\pi^{0}}$ in AllPar, cf. Table 5). Overall, this leads to a reduction of $75.31$ in $\chi^{2}$ (Table 5), similar to G24-c, as well as a good data-MC agreement across all data sets.
Refer to captionRefer to captionRefer to caption
V Summary and Outlook
@@
This work represents the first global tuning effort on TKI data. Our partial tune of the SF-LFG and hA models, G24_20i_06_22c (G24-c), provides an effective theory to better describe both the neutrino-hydrocarbon pionless and pion production. The largest change in the model is demanded by the MINERvA-$\pi^{0}$ TKI measurement [12] that was significantly overestimated in genie. The improvement is crucial for future accelerator-based neutrino experiments. This tuning configuration has been integrated into the master branch of genie and is slated for inclusion in the upcoming release.
@@
To develop an effective theory, we focused on the most sensitive parameters, $R_{\textrm{SRC}}$, $S_{\lambda}^{\pi^{0}}$, $S_{\textrm{CEX}}^{\pi}$, $S_{\textrm{CEX}}^{\textrm{N}}$, $S_{\textrm{ABS}}^{\textrm{N}}$, and $S_{\textrm{PIPD}}^{\textrm{N}}$, of the SF-LFG and hA models, reducing the total number of parameters from $14$ to $6$. The corresponding best combination of observables is $\delta\alpha_{\textrm{T}}$, $p_{\textrm{N}}$ (or $\delta p_{\textrm{T}}$ if no $p_{\textrm{N}}$ is available), and $\delta p_{\textrm{TT}}$, from pionless and pion-production measurements in both T2K and MINERvA. In doing so, the pion production model is held fixed such that the tuned values of $S_{\lambda}^{\pi^{0}}$, $S_{\textrm{CEX}}^{\pi}$, $S_{\textrm{ABS}}^{\textrm{N}}$, and $S_{\textrm{PIPD}}^{\textrm{N}}$ depart from the established constraints from hadron measurements. A different pion production model might compensate part of these model tuning effects. We also chose the hA FSI model for practical reasons. Its sequential treatment of MFP and rescattering should be considered as a numerical advantage (as was also pointed out in Ref. [9]). Other more sophisticated models could be considered in future tunes. With the existing four TKI measurements from T2K and MINERvA, we have also derived other effective tunes like G24_20i_fullpara_alt (G24-t). The degeneracy can be resolved with additional data, particularly from argon-based measurements [20, 21, 22, 23, 24, 61].
@@ Acknowledgment
We would like to thank the CC-IN2P3 Computing Center providing computing resources and for their support. X.L. is supported by the STFC (UK) Grant No. ST/S003533/2. J.T.-V. acknowledges the support of the Raymond and Beverly Sackler scholarship.
appendix
.1 Combinations of Observables
@@
.2 $\pi^{0}$ FSI Fate Decomposition for G24-0 and G24-c
Figures 14-16 display comparisons of G24-0 and G24-c predictions to data, detailing the composition by $\pi^{0}$ FSI fate.
Combi- 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 (Best- AllPar) 16 17 18 19 20 21 22 23 24 (Super- set) 25 26 (Best- RedPar)
T2K-$0\pi$
T2K-$\pi^{+}$
MINERvA-$0\pi$
MINERvA-$\pi^{0}$
Table 6: Specifications of observable combinations within the tuning superset in Table 3.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 13: Extension of Fig. 10a to all four $\delta\alpha_{\textrm{T}}$ measurements.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 14: Extension of Fig. 10a but to all four $p_{\textrm{N}}$ measurements.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 15: Extension of Fig. 10b to all four $\delta\alpha_{\textrm{T}}$ measurements.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 16: Extension of Fig. 10b but to all four $p_{\textrm{N}}$ measurements.

### v2

@@ 5.2 Alternative tune: G24-t
A similar improvement is achieved by the intermediate tune, G24-t, illustrated in Fig. 12. Instead of significantly increasing proton PIPD to elevate the $p_{\textrm{N}}$ tail, this tune notably enhances $R_{\textrm{SRC}}$—less-energetic final states are present and therefore an increased RES/DIS (Fig. 12a) ratio. Rather than suppressing $S_{\textrm{CEX}}^{\pi}$ (cf. Fig. 17d in Appendix B), this tune significantly increases it (Fig. 12b) to offset the further reduction of “No Interaction” $\pi^{0}$ events (smaller $S_{\lambda}^{\pi^{0}}$ in AllPar, cf. Table 5). Overall, this leads to a reduction of $75.31$ in $\chi^{2}$ (Table 5), similar to G24-c, as well as a good data-MC agreement across all data sets.
Figure 13: Change in MC prediction for $\pi^{0}$ cross section between G24-0 and G24-c.
5.3 Discussion
For both G24-c and G24-t the fit results suggest extreme values, but the effect of the parameter change is less effective than the change in the parameter suggests. For the individual FSI fate cross section, due to the re-normalization step after scaling, the effective change for one particular fate is smaller than the scaling parameter magnitude as explained in Sec. 4.2. As for the total FSI cross sections, only that of $\pi^{0}$ is changed. The total $\pi^{0}$-C scattering cross section for pion kinetic energy is plotted in Fig. 13 to assess the overall impact. The shape remains similar and consistent with Fig. 2.23 ($\pi^{+}$-C reactions) in Ref. Andreopoulos:2015wxa while the peak magnitude increases from $380$ to $538$ mb, a $48\%$ increase, considerably less than the scaling factor, $S_{\lambda}^{\pi^{0}}=0.22$, seemingly suggests. Moreover, the default $\pi^{0}$ parameter values are calculated from charged pion data assuming isospin symmetry rather than extracted directly from experimental data. Hence, this modification does not violate any existing agreement with hadron scattering data.
In addition, agreement of this tune with most of the non-TKI neutrino datasets available is unchanged before and after the tune, thereby demonstrating the physicality of the tune. In conclusion, this tune is a valid and effective model that can be used as a starting point for an analysis. Further refinement will be available once more data will be included in the fit, including non-TKI observables.
6 Summary and outlook
@@
This work represents the first global tuning effort on TKI data. Our partial tune of the SF-CFG and hA models, G24_20i_06_22c (G24-c), provides an effective theory to better describe both the neutrino-hydrocarbon pionless and pion production. The largest change in the model is demanded by the MINERvA-$\pi^{0}$ TKI measurement MINERvA:2020anu that was significantly overestimated in genie. The improvement is crucial for future precision GeV neutrino experiments. This tuning configuration has been integrated into the master branch of genie and is slated for inclusion in the upcoming release.
@@
To develop an effective model, we focused on the most sensitive parameters, $R_{\textrm{SRC}}$, $S_{\lambda}^{\pi^{0}}$, $S_{\textrm{CEX}}^{\pi}$, $S_{\textrm{CEX}}^{\textrm{N}}$, $S_{\textrm{ABS}}^{\textrm{N}}$, and $S_{\textrm{PIPD}}^{\textrm{N}}$, of the SF-CFG and hA models, reducing the total number of parameters from $14$ to $6$. The corresponding best combination of observables is $\delta\alpha_{\textrm{T}}$, $p_{\textrm{N}}$ (or $\delta p_{\textrm{T}}$ if no $p_{\textrm{N}}$ is available), and $\delta p_{\textrm{TT}}$, from pionless and pion-production measurements in both T2K and MINERvA. In doing so, the pion production model is held fixed such that the tuned values of $S_{\lambda}^{\pi^{0}}$, $S_{\textrm{CEX}}^{\pi}$, $S_{\textrm{ABS}}^{\textrm{N}}$, and $S_{\textrm{PIPD}}^{\textrm{N}}$ depart from the established constraints from hadron measurements. A different pion production model might compensate part of these model tuning effects Yan:2024kkg . We also chose the hA FSI model for practical reasons. Its sequential treatment of MFP and rescattering should be considered as a numerical advantage (as was also pointed out in Ref. GENIE:2022qrc ). Other more sophisticated models could be considered in future tunes. With the existing four TKI measurements from T2K and MINERvA, we have also derived other effective tunes like G24_20i_fullpara_alt (G24-t). The degeneracy can be resolved with additional data, particularly from argon-based measurements MicroBooNE:2022emb ; MicroBooNE:2023cmw ; MicroBooNE:2023tzj ; MicroBooNE:2023wzy ; MicroBooNE:2024tmp ; MicroBooNE:2015bmn .
@@ Acknowledgment
We thank the CC-IN2P3 Computing Center providing computing resources and for their support. X.L. is supported by the STFC (UK) Grant No. ST/S003533/2. J.T.-V. acknowledges the support of the Raymond and Beverly Sackler scholarship.
Appendix A Combinations of observables
Combi- 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26
T2K-$0\pi$
T2K-$\pi^{+}$
MINERvA-$0\pi$
MINERvA-$\pi^{0}$
Table 6: Specifications of observable combinations within the tuning superset in Table 3. Combi-15 is Best-AllPar, Combi-24 is Superset, and Combi-26 is Best-RedPar.
@@ Appendix A Combinations of observables
Appendix B $\pi^{0}$ FSI fate decomposition for G24-0 and G24-c
Figures 15-17 display comparisons of G24-0 and G24-c predictions to data, detailing the composition by $\pi^{0}$ FSI fate.
Figure 14: Extension of figure 10a to all four $\delta\alpha_{\textrm{T}}$ measurements.
Figure 15: Extension of figure 10a but to all four $p_{\textrm{N}}$ measurements.
Figure 16: Extension of figure 10b to all four $\delta\alpha_{\textrm{T}}$ measurements.
Figure 17: Extension of figure 10b but to all four $p_{\textrm{N}}$ measurements.